A 3.9−μF capacitor is connected to a 8.8−V battery. What is the magnitude of the charge on each plate?

Q = μC.
Answer #1

Then C=Q/V

SO,

Q=CV

Q=3.9*10^-12*8.8

Q=3.43*10^-11 C

Q=3.43 uC Answer

I hope help you !!